OpenStudy (anonymous):

What is the equation of the line that passes through the point (0, 3) and is perpendicular to the line y = 2x + 3?

OpenStudy (cwrw238):

perpendicalar means that the line's slope = -1 / slope of y = 2x + 3 (has slope 2) so itsslope = - 1/2 use the standard formula y - y1 = m(x - x1) where m = slope and x1, y1 is a point on line y - 3 = -1/2 ( x - 0)

OpenStudy (cwrw238):

can you simplify that last equation?